There is a long line of eager children outside of your house for trick-or-treating, and with good reason! Word has gotten around that you will give out 3k pieces of candy to the kth trick-or-treater to arrive. Children love you, dentists despise you.

(a) Expressed in summation notation (using a Σ), what is cn, the total amount of candy that you should buy to accommodate n children total?

(b) Use induction to prove that the total amount of candy that you need is given by the closed-form solution: cn = (3n+1 - 3) / 2 


1
Expert's answer
2021-10-26T16:06:52-0400

a)

"c_n=\displaystyle{\sum_{k=1}^n}3^k"


b)

for k=1

"c_1=3=\frac{3^2-3}{3}"


let

"c_n=\displaystyle{\sum_{k=1}^n}3^k=\frac{3^{n+1}-3}{2}"


then:

"c_{n+1}=\displaystyle{\sum_{k=1}^{n+1}}3^k=\displaystyle{\sum_{k=1}^n}3^k+3^{n+1}=\frac{3^{n+1}-3}{2}+3^{n+1}=\frac{3^{n+1}-3+2\cdot3^{n+1}}{2}="


"=\frac{3\cdot3^{n+1}-3}{2}=\frac{3^{n+2}-3}{2}"